OpenStudy (anonymous):

Solve S = 3f – 12 for f.

jimthompson5910 (jim_thompson5910):

\[\Large S = 3f – 12\] \[\Large S+12 = 3f\] \[\Large \frac{S+12}{3} = f\] \[\Large f=\frac{S+12}{3}\]